A magnetised wire of moment 3.14 A-m^2 is bent in the form of a semicircle, then the new magnetic moment will be:-
  1. 3.14 Am^2
  2. 2
  3. 6.28
  4. none
explain your answer

Grade:

1 Answers

Khimraj
6 years ago
Let pole strength of wire is p and length is L , then
p*L = 3.14 Am2
now it bend in semicircle
L = 3.14R.
R = L/3.14
effective length of semicircle = 2R = 2L/3.14
New magnetic moment of wire = p*(2R) = (2/3.14)*p*L = (2/3.14)*(3.14) = 2.
Hope it clears If you likeanswer then please approve it.